# All-in-one: Understanding and Generation in Multimodal Reasoning with the MAIA Benchmark

### v2

@@ Acknowledgments
This work has been carried out while Davide Testa was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with Fondazione Bruno Kessler (FBK). Giovanni Bonetta and Bernardo Magnini were supported by the PNRR MUR project PE0000013-FAIR (Spoke 2). Alessandro Lenci was supported by the PNRR MUR project PE0000013-FAIR (Spoke 1). Alessio Miaschi was supported by the PNRR MUR project PE0000013-FAIR (Spoke 5). Lucia Passaro was supported by the EU EIC project EMERGE (Grant No. 101070918). Alessandro Bondielli was supported by the Italian Ministry of University and Research (MUR) in the framework of the PON 2014-2021 “Research and Innovation" resources – Innovation Action - DM MUR 1062/2021 - Title of the Research: “Modelli semantici multimodali per l’industria 4.0 e le digital humanities.”
Appendix A MAIA: Benchmark Creation Process
This appendix provides a detailed explanation of the step-by-step process involved in developing the dataset of the benchmark. Each sub-section corresponds to a specific stage, outlining the methodologies, decisions, and criteria that guided the creation of MAIA as an evaluation framework. Figure 4 illustrates the underlying logic of the evaluation framework, outlining its construction process and the tasks involved. Similarly, Table 5 shows MAIA’s dataset structure after the creation and validation process which is accurately described in the following Section. The final structure of the dataset is shown in Figure 5, which accurately provides examples of aligned question, answer, TS, and FS.
Figure 4: Workflow of the MAIA evaluation framework, integrating dataset construction with its application to the two main benchmark tasks.
Feature n
Videos $100$
Semantic Categories 1212121211footnotemark: 1 ($9$ Macro-Cat.)
Questions $2,400$ ($2$ x Category)
8-Answers Pool $2,400$
Answers $19,200$
True Statements $19,200$
False Statements $19,200$
Table 5: MAIA Statistics
Figure 5: Overview of MAIA reasoning categories. For each of the 100 videos, it contains 2 questions for each of the 12 categories; for each question, it has 8 answers, and each of these answers has a corresponding True and False statement pair.
A.1 Video Selection
The first step in creating the benchmark involved the selection of a culturally representative dataset of $100$ short videos coming from YouTube, each lasting approximately $30$ seconds. To reflect Italian culture authentically and to ensure richness and variety, the selection process tried to be focused on the following thematic areas:
  • Locations
  • Food
  • Sport
  • Job
  • Nature
  • Activities
Such topics allowed us to collect a dataset showing locations (indoor and outdoor), iconic Italian cities, and daily activities such as crossing the street, enjoying breakfast at a café, cooking pasta, attending a soccer match or even important events such as Italian local festivals or typical Italian weddings. Priority was given to videos featuring people, particularly with close-up shots, to enhance visual and contextual relevance.
The video collection process was facilitated by an automated script that scraped videos from YouTube matching specific keywords tied to these thematic areas. Additional filters ensured that the content came from YouTube Italy and respected the Creative Commons licensing for the content use and reproducibility. Once the videos were identified, as many of them were too long, the most engaging $30$-second segments were extracted and cut to create a uniform set of $100$ short clips, representing the base of our Italian-native dataset.
A.2 Semantic Categories Definition
This step focused on establishing and refining the semantic categories and phenomena that would form the foundation for analyzing the previously selected videos. This process aimed to ensure that the benchmark comprehensively addressed a wide range of linguistic reasoning categories, providing a robust framework for evaluating vision-language models. Each of these categories was carefully tailored to align with the cultural and contextual richness of the selected video dataset and resulted from a long period of discussion and refinements. Initial categories were proposed and systematically tested against sample videos to assess their applicability and coverage. Feedback from these tests was used to refine definitions, ensuring clarity and reducing overlap between categories. The final set of semantic categories represents a broad spectrum of linguistic and semantic reasoning phenomena, designed to challenge VLMs in interpreting the intricate interaction between visual and linguistic cues in diverse, real-world contexts. These categories form the backbone of the benchmark, allowing the models’ reasoning capabilities to be extensively tested on culturally representative Italian scenes. Next, the following paragraphs present in detail the nine final macro-categories, along with their definitions and some examples of usage. These categories and descriptions were also used to administer the Google forms for collecting data from the annotators.
Causal
@@ A.2 Semantic Categories Definition
A.3 Data collection
This section outlines the data collection process and the methodologies adopted for this step of the research which plays a crucial role in the development of this evaluation framework. This section is divided into two main parts: Questions Collection, which describe the procedure for creating relevant questions based on video content, and Answers Collection, which tells how annotators provided diverse responses to those questions.
Questions Collection
After defining the Semantic Categories on which we based the subsequent analysis, we created $12$ different sets of guidelines (one for each subcategory [^11] The original subcategories were $13$, but due to the difficulty of the causal category, we decided to merge the category of explicit and implicit causals into one group. The category was then assigned to a student who annotated the first $50$ videos with explicit causals and the last $50$ with implicit causals to allow the subcategories to be balanced within the same group.) that were administered to the same number of annotators via Google Forms.
Figure 6: Example of a Google Form used for collecting 2 Questions for each video for each of the assigned category
All the selected annotators were PhD students under $30$ with specializations in Linguistics and Computational Linguistics. They were paid with $100$ euros for the generation task, which was executed in multiple stages through the administration of $1,200$ different forms ($10$ forms per annotator). After carefully reading the rules and instructions in the form, their task was to watch the video as much as they wanted, and to compulsorily generate $2$ questions (i.e., Question_A and Question_B) for each of the $100$ videos in the dataset, as shown in Figure 6 . To ensure variability between the pair of questions about that video, in the second question, the annotators were asked to try as far as possible to change the entities and/or events involved in the first one previously formulated. The final result corresponds to a total of $2,400$ questions ($24$ per video). Figure 7 shows an example of a video with $12$ out of $24$ questions for each semantic category.
Figure 7: Example of a video with its $12$ Questions, each one for a sub-category.
Each form provided contained both the definition of the assigned semantic category with examples (See section A.2), and also general rules to be followed. Each question had to be generated naturally and as an open-ended question. Questions involving a ‘Yes/No’ answer (e.g. ‘Is there a car in the video?’) were not allowed. Finally, for the correct execution of the task, the audio of the video had to be ignored, as the Vision Language Models to be tested could only work on the visual part.
Subsequently, these questions were manually reviewed to ensure quality and category alignment. Thus, while verifying whether the questions generated could actually belong to the specific category and whether they complied with the rules presented in the forms, one third of the generated questions were modified or refined.
Answers Collection
Once the questions were obtained and checked, they were used to set up the answer collection process through Prolific, a platform designed to help researchers recruit participants for high-quality data collection. To follow the initial goal of creating an Italian-native resource that accurately represents Italian culture, we implemented a preliminary screening process to ensure that participants would provide answers from the point of view of an Italian person. For this data collection step, we accepted only annotators aged $25$ to $80$ who were born in Italy, spoke Italian as their first language, and had spent the majority of their first 18 years of life in Italy.
As with the question collection step, we used Google Forms (Figure 8) to provide the task to the annotators that was paid £$7$ per hour. Each form included $10$ videos, and for each video, the annotators were asked to answer $12$ questions [^12] Some annotators were given the $12$ Questions_A for that video and others the $12$-Questions_B group (one per category). The forms provided detailed guidelines on how to answer and specified the criteria for an acceptable response, which was necessary for the annotators to receive payment. Each response needed to be a complete answer (i.e., not short answers with missing noun / prepositional phrases), rely only on visible or deducible information from the video, and exclude any information derived from the video’s audio. Annotators were also encouraged to use their own world knowledge when interpreting the visual content.
Figure 8: Example of a Google Form used for collecting answers for each video
The initial acceptability of the task was determined by electing $2$ control questions from the $12$ semantic categories: the Out-of-scope and the Uncertainty question. Both questions had predefined acceptable answers, which made them easy to use for this purpose: the former required the negation of what asked in the question, while the latter required an expression of uncertainty in the response. To evaluate the acceptability of these control question answers, we employed the GPT-4o model in a few-shot prompting setting. During each annotation task, annotators worked with $10$ videos, each containing the $12$ questions,including the control questions. Annotation work was accepted only if the annotators correctly answered at least $90\%$ of the control questions (i.e., $18$ out of $20$ correct answers); otherwise, their submissions were rejected, and the task was reassigned to another annotator.
Our goal was to collect 8 different answers for each question to ensure not only accuracy but also variability in responses. To achieve this, we presented the same question to $8$ different annotators, collecting their individual answers. In total, $2,400$ questions were paired with 8 answers each, resulting in $19,200$ responses.
These responses then were features by a semi-automated review and validation process divided into two main steps, both facilitated by GPT-$4$o with few-shot prompting.
  1. 1.
    Semantic Consistency check. Each response was evaluated for semantic consistency with the corresponding question (Figure 9A). In cases where inconsistencies were detected, the responses were manually reviewed to determine whether the question should be re-answered by another annotator or the response could still be accepted. Real inconsistencies were found to be minimal (i.e., fewer than $100$ out of $19,200$ responses).
  2. 2.
    Contradiction test. By using GPT-4o OpenAI et al. (2024), we checked whether, within each pool of $8$ responses to the same question, any of the response contradicted the others (Figure 9B). We found that 90.25% of the 8-answer pools exhibit full agreement, as they do not contain any contradictions. The remaining 9.75% (234 cases) was manually reviewed by an additional annotator to resolve inconsistencies. The proportion of non-contradiction within the human annotated answers can also be considered as a rough proxy for Inter-annotator Agreement (IAA), since traditional IAA metrics are not easily applicable due to the open-ended nature of MAIA questions.
Figure 9: Prompts used for Answers’ semi-automatic validation. Prompts were provided with some examples (i.e., few-shot prompting) but here in the Figure only the main instructions are presented.
After completing these reviews, we obtained a dataset of $2,400$ questions, each paired with a pool of $8$ high-quality responses, resulting in a total of $19,200$ validated responses.
A.4 Answers Post-Processing
Since the data obtained up to this point will form the solid basis for the subsequent benchmarking phases (A.5 and A.7), we decided to set up a post-processing phase to assess and check that a certain degree of variability was ensured within each of the $2,400$ $8$-Answers pool. This work was carried out in several incremental stages, each of which involved an initial analysis of the potential redundancy of responses within the pool. Each of these analyses was then followed by a specific phase of semi-automatic expansion of these responses through the introduction of lexical variability.
  1. 1.
    Equal strings: a classic check to see if there were any pairs of equal answers. Table 6 shows the number of identified equal pairs divided by semantic category, highlighting a higher score for Out-of-scope Answers. Based on these results, GPT-4o was used in a Sentence Rephrasing task in a zero-shot prompting setting (Figure 10A). The only exception was the Out-of-scope category where all 1600 responses were manually edited.
  2. 2.
    Lexical Overlap: we tokenized each answer, checked the number of Type-Tokens shared between answer pairs within the pool of 8 and selected those that shared more than $50$% of them. This phase was followed for these selected items by an automatic Sentence Rephrasing task in which GPT-4o generated a new Answer with a Lexical overlap < $50$% with reference to the other $7$ Answers in the pool (Figure 10B). Table 7 shows the lexical overlap statistics before and after reformulation. The final average lexical overlap score within a pool($18.74$) represents the most balanced situation achievable. Indeed, it was noted that further attempts at lexical modifications within the pool would increase this percentage rather than make it lower. Finally, by focusing only on tokens represented by content words (i.e. adjectives, nouns, verbs and adverbs), it was noted that the lexical overlap index drops even further to $17.60$%.
Figure 10: Prompts used for Answers’ post-processing with GPT4-o
These two checks were finally followed by an analysis performed using the Type-Token Ratio (TTR) as a measure to quantify lexical richness (Table 7). The final results showed that, considering only content words, only $633$/$2400$ pools had a TTR < $0.5$ and that the average TTR per pool stands at $0.55$.
Category Equal Pairs
Causal 21
Counterfactual 2
Partial Implicit 133
Total Implicit 93
Uncertainty 112
Out-of-scope 414
Planning 2
Sentiment 9
Partial Spatial 77
Total Spatial 101
Temporal Duration 91
Partial Temporal 9
Tot. 1064
(2128 / 19200 Answers)
Table 6: Number of equal answers pairs in our
dataset before post-processing divided by semantic categories.
Before Rephrasing After Rephrasing
CW-only
Pool Avg. Lexical Overlap $22.95$% $18.74$% $17.60$%
Pool with Lexical Overlap > $50$% $11$ 00 00
Avg TTR for pool $0.50$ $0.55$
Pool with TTR < $50$% (Avg.) $1197$/$2400$ $633$/$2400$
Table 7: Lexical Overlap and TTR Statistics (Type-token) before and after Automatic Sentence Rephrasing with GPT-4o for $19,200$ Answers. Content-words (CW only) statistics are shown only for the After Rephrasing part.
After these steps, Figure 11 represents a valuable example of a high-quality question with a pool of $8$ answers in which we can find a significant level of lexical variability.
Figure 11: Example of a video with a Questions (Implicit category) and its 8-Answers pool.
A.5 True Statement Generation
After the collection and validation of the questions and answers, the next step involved the automatic generation of true statements (TS) which consists of descriptive declarative sentences aligned with the visual content of the videos. Just as a caption describes an image, a TS describes a video (or part of it) from different perspectives according to the specific semantic category.
For example, suppose a video shows a boy who is initially in a kitchen and while he is cooking he hears a loud noise and runs away to another room. An example of TS for the Spatial category could be:
Italian: Nel video il ragazzo è in cucina prima di scappare via.
English: In the video the boy is in the kitchen before running away.
In contrast a Causal TS:
In the scene the boy runs away because of the loud noise.
To create such statements, we used, as in the previous steps, GPT-4o (see Figure 12A for the prompt used). Using the information conveyed by the single question and its answer, we were able to combine them, generating $19,200$ true statements (TSs). As with the answers, the TS are also organised into $2,400$ pools of $8$ items. Each of these (TSs) in a pool will express the same event with different words and/or points of view.
A.6 True Statement post-processing
Just like with the pools of 8 responses, we also performed checks and semi-automatic interventions to ensure lexical variability within the 2400 pools of true statements (TS).
The steps followed were the same, except that for the TS we used the Type-Token Ratio (TTR) as an additional tool to enhance lexical variability within the pools. First, we identified $326$ pairs of identical TS and addressed this by regenerating the sentences using GPT-4o, as already done in LABEL:subsec:A-postprocessing, with the prompt shown in Figure 10A. Then, we analyzed the lexical overlap within the pools, focusing on statements with an overlap greater than $50$% ($354$ cases). This reduced the average overlap percentage of a pool from $39.34$% to $30.51$% (Type-Token) as shown in Table 8.
As a final step, we used the TTR to increase lexical richness within individual pools. We identified pools with a TTR < $0.5$ and replaced content words with synonyms to ensure that each TS contained different words. This approach allowed us to achieve a TTR of $0.50$ ($0.55$ when considering only content words), which is an excellent result given that, unlike responses, TS have a more fixed and repetitive structure (e.g., "In the video X happens" "The video shows X" etc.).
Before Rephrasing After Rephrasing
Type-Token CW-only Type-Token CW-only
Pool Avg. Lexical Overlap $39.34$% $38.04$% $30.51$% $26.81$%
Pool with Lexical Overlap > $50$% $354$ $344$ 00 00
Avg TTR for pool $0.37$ $0.41$ $0.50$ $0.55$
Pool with TTR < $50$% (Avg.) $2288$ $1964$ $1196$ $634$
Table 8: Lexical Overlap and TTR Statistics (Type-token and Content word-only) before and after Automatic Sentence Rephrasing with GPT-4o for 19200 TSs.
A.7 False Statement Generation
The automatic generation of true statements represents the final phase of data creation, ensuring all elements are ready to begin the actual benchmark task. Starting from the notion of the dichotomy of Caption and Foil —where a caption is a brief and correct descriptive sentence about a scene, and a foil is a minimally altered version of the caption that turns it into an incorrect description— we developed our False Statements (FSs).
The final goal of this work is to create minimal pairs for each semantic category, enabling controlled experiments and precise analysis of a model’s behavior with respect to the specific category.
As for TSs, the FSs were also automatically generated using $GPT-4o$. Figure 12B illustrates the prompt used during this phase, where the task was presented to the model using the Caption/Foil terminology rather than True/False Statements to enhance the effectiveness of the output. Given a semantic category and a caption, the task was to generate the corresponding foil by editing only the elements of the sentence related to that semantic category (e.g., If the category is Spatial, the prepositional phrase conveying spatial or location information is the block to be edited for creating the corrisponding foil).
@@ A.7 False Statement Generation
A.8 False Statement Validation
As the last step of this work phase concerning the generation of data, we focused on the validation of the false statements we had just obtained. This review was characterised by the presence of two different semi-automatic checks based on the actual definition of foil: A description of the visual content minimally different from the caption, such that this difference falsifies its veracity. In both cases, GPT-4o was used as a model for the implementation of the task in a semi-automatic fashion.
  • Structural check: a minimal prompt (see Figure 13A) was constructed to check the correctness of a foil with respect to the related caption according to the specific category (e.g., if the category is Spatial, investigate whether, given a caption and a foil, the foil is correct with respect to the spatial information). Overall, $2.5$% of the foils generated turned out to be incorrect and then manually fixed. Table 9 shows for each semantic category the number of foils that resulted as incorrect according to GPT-4o, as opposed to those that actually were not and that were then manually corrected.
  • Contradiction test: based on the notion that the correct foil must be in contradiction with the relevant caption, we then ran an NLI task to identify cases of Entailment, Neutrality and Contradiction between our pairs of True and False Statements(Figure 13B). After a qualitative analysis of these cases, we realised that the cases identified by the model as Neutral (1287), were in fact correct and already partly defined a contradiction as in the next example.
    Italian:
    TS: I bambini toccano la lana per capirne la consistenza.
    FS: I bambini toccano la lana per capirne il colore.
    English:
    TS: The children touch the wool to check its texture.
    FS: The children touch the wool to check its colour.
    For this reason, the focus shifted only to the cases defined as Entailment ($170$). Within this group, only $93$ turned out to be true Entailments and they were subsequently manually modified.
This last revision allow us to obtain a packet of $19,200$ high-quality FFs with respect to its relative TSs.
Semantic Category FSs Incorrect FSs
GPT-4o Human
Causal 1600 80 54
Counterfactual 1600 61 15
Partial Implicit 1600 94 38
Total Implicit 1600 99 63
Uncertainty 1600 27 20
Out-of-scope 1600 48 9
Planning 1600 159 117
Sentiment 1600 10 7
Partial Spatial 1600 149 58
Total Spatial 1600 33 14
Temporal Duration 1600 33 29
Partial Temporal 1600 71 55
Tot. 19200 864 479
Table 9: Results of Semi-automatic False Statements Validation (Structural check)
Figure 13: Prompts used for False Statement Validation: Structural check (A) and Contradition Test (B)
@@ Appendix B Experiments
This appendix section will contain additional details on our experimental settings, including a description of the Vision-Language Models used, as well as graphs and tables summarizing the results for Tasks 1 and 2 of MAIA. These supplementary materials provide a more comprehensive view of our findings, which could not be included in the main paper due to space limitations.
Note that in contrast to the initial experiments for creating and validating the synthetic data of the MAIA dataset, where we used OpenAI’s GPT-4o API, the experiments on MAIA were conducted using A100100100100 GPUs ($40$GB). Overall, the total computational budget was on the order of \$̃1,000\$ GPU hours.
@@ B.1 Models tested
We benchmarked four recent VLMs, sourced from the Hugging Face hub, representing state-of-the-art approaches in Vision-Language tasks. Both experiments and related quantitative evaluation have been done using lmms-eval (Zhang et al., 2024a), a framework for the evaluation of multimodal models.
InternVL2 Chen et al. (2024): $8$B parameter transformer-based multimodal model employing advanced cross-attention; pre-trained on large-scale image-text and video datasets for diverse multimodal tasks and instruction-tuned. It uses InternLM2.5 as open-sourced-$7$B parameter chat model.
Hugging Face model: OpenGVLab/InternVL2-8B
LLaVA-NeXT-Video Zhang et al. (2024b): $7$B parameter model built on the LLaVA framework, optimized for video understanding with mechanisms to capture temporal dynamics; fine-tuned on video instruction data. Base LLM: Vicuna-$7$B (v1.51.51.51.5).
Hugging Face model: llava-hf/LLaVA-NeXT-Video-7B-hf
LLaVa-OneVision Li et al. (2024a): $7$B parameter model that builds on the LLaVA framework with a Qwen2 LLM backbone to serve as a general-purpose vision-language assistant; pre-trained on extensive multimodal data to deliver robust cross-modal reasoning.
Hugging Face model: lmms-lab/llava-onevision-qwen2-7b-ov
Qwen2.5-VL Qwen et al. (2025): the latest $7$B parameter VLM of the Qwen family using the Qwen2.52.52.52.5 LLM decoder; key enhancements are related to grounding, working with longer videos and capturing events. It was pre-trained on comprehensive visual and textual datasets and fine-tuned for detailed, context-aware responses.
Hugging Face model: Qwen/Qwen2.5-VL-7B-Instruct
@@
As described in Section 4, we used MAIA for benchmarking also some LLMs in a text-only evaluation. We used five open-weight LLMs which have shown good performance on a variety of tasks on Italian Magnini et al. (2025). For conducting these experiments, we used Minicons library Misra (2022), a high-level wrapper around Hugging Face for investigating predictive behavior of transformer models. Specifically, probabilities were computed adding a normalization parameter to take into account the different length of sentences in terms of tokens. Models used in the Hugging face Hub are: Llama-$3.1$ ($8$B-Instruct), LLaMAntino-$2$ ($7$B), LLaMAntino-$3$-ANITA ($8$B-Instruct), Gemma ($7$B) and Qwen2.52.52.52.5 ($7$B-Instruct).
B.2 Task1: Visual Statement Verification
As described in Section 4.4, the performance of the models in this task was compared against our most probable baseline. Figure 14 illustrates this relationship for each semantic category. Interestingly, text-only LLMs sometimes achieve higher accuracy in selecting the correct statement than VLMs, despite lacking visual input. For instance, in the counterfactual category the baseline values closely match those of the VLMs, while in the uncertainty category the baseline even outperforms both InternVL and LLaVA-NeXT-Video. Overall, the accuracy gap between our baseline and the VLMs is generally smaller than one might expect or theoretically anticipate. Table 10 presents the results of Task $1$, providing a detailed analysis for the Causal category, distinguishing between Explicit and Implicit. The same analysis is conducted in Table 13 for Task $2$ taking into account LLM’s judgements. Figure 15 represetns models’ fingerprint through MAIA’s reasoning categories, specifically illustrating models’ performance in the NLU Task 1 while Figure 16 reports performance on the same task with MAIA’s 8-answer pools.
Explicit Implicit Partial Total Partial Total Duration Partial
Unimodal 0.56 0.43 0.48 0.73 0.55 0.68 0.53 0.79 0.50 0.51 0.48 0.51 0.53 0.48
Black video InternVL2 0.68 0.61 0.67 0.88 0.89 0.69 0.61 0.96 0.52 0.59 0.54 0.55 0.67 0.60
Llava-Next-Video 0.50 0.48 0.50 0.52 0.48 0.51 0.52 0.45 0.50 0.51 0.51 0.49 0.51 0.48
Llava-oneVision 0.59 0.52 0.57 0.48 0.92 0.61 0.38 0.97 0.52 0.53 0.50 0.52 0.60 0.51
Qwen-2.5-VL 0.73 0.66 0.70 0.86 0.87 0.75 0.72 0.98 0.56 0.62 0.61 0.65 0.73 0.70
1-Frame InternVL2 0.75 0.77 0.82 0.87 0.69 0.71 0.77 0.89 0.65 0.80 0.65 0.82 0.67 0.69
Llava-Next-Video 0.61 0.67 0.70 0.76 0.49 0.63 0.75 0.40 0.59 0.65 0.59 0.70 0.56 0.53
Llava-oneVision 0.76 0.75 0.83 0.78 0.89 0.74 0.76 0.91 0.68 0.81 0.64 0.81 0.70 0.67
Qwen-2.5-VL 0.79 0.79 0.83 0.81 0.91 0.75 0.82 0.70 0.82 0.96 0.67 0.82 0.69 0.73
32-Frames InternVL2 0.79 0.82 0.84 0.85 0.77 0.75 0.84 0.84 0.75 0.83 0.69 0.86 0.66 0.76
Llava-Next-Video 0.52 0.54 0.57 0.57 0.42 0.57 0.61 0.32 0.52 0.59 0.52 0.56 0.51 0.48
Llava-oneVision 0.81 0.86 0.88 0.78 0.88 0.76 0.88 0.85 0.80 0.85 0.71 0.87 0.65 0.80
Qwen-2.5-VL 0.84 0.89 0.90 0.80 0.89 0.78 0.86 0.92 0.82 0.88 0.83 0.90 0.75 0.81
Table 10: Visual statement verification (Task $1$): accuracy of correct choices across reasoning categories with Causal sub-categories specification.
Figure 14: Accuracy of 4 VLMs compared with our most-probable baseline with reference to each category in MAIA
Figure 15: Task 1
Figure 16: Task 1 Aggregate
Models consistency.
In Task 1, model consistency in making uniform decisions within the same pool of True-False Statement pairs was investigated through a quantitative analysis, as outlined in Section 5.1. Moreover, Table 12 reports the consistency of models across the same pool of TF-FS pairs in Task $1$. As we can see, imposing a fair and rigid evaluation regime, viz. evaluating a model positively only when consistent across datapoints (total agreement), the 1-frame bias is mitigated for almost all models. The most consistent model is Qwen2.5-VL, while LLaVA-NeXT-Video performs best when a simple majority (at least $4$/$8$ consistent decisions) is considered.

### v3

@@ Acknowledgments
This work has been carried out while Davide Testa was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with Fondazione Bruno Kessler (FBK). Giovanni Bonetta and Bernardo Magnini were supported by the PNRR MUR project PE0000013-FAIR (Spoke 2). Alessandro Lenci was supported by the PNRR MUR project PE0000013-FAIR (Spoke 1). Alessio Miaschi was supported by the PNRR MUR project PE0000013-FAIR (Spoke 5). Lucia Passaro was supported by the EU EIC project EMERGE (Grant No. 101070918). Alessandro Bondielli was supported by the PNRR MUR project PE0000013-FAIR (Spoke 1), funded by the European Commission under the NextGeneration EU programme and by the the Italian Ministry of University and Research (MUR) in the framework of the PON 2014-2021 “Research and Innovation" resources – Innovation Action - DM MUR 1062/2021 - Title of the Research: “Modelli semantici multimodali per l’industria 4.0 e le digital humanities.”
Appendix A MAIA:Benchmark Details
A.1 Semantic Categories Definition
We report here the definition of the twelve reasoning categories included in MAIA.
CAUSAL.
@@ A.1 Semantic Categories Definition
A.2 True and False Statement Generation
Figure 4 illustrates the prompts used to generate True Statements (A) from the questions and the corresponding responses in the 8-answer pools and the False Statements (B) starting from the true ones.
@@ Appendix B Experiments
This appendix section will contain additional details on our experimental settings, including a description of the VLMs used, as well as graphs and tables summarizing the results for Tasks 1 and 2 of MAIA.
In contrast to the initial experiments for creating and validating the synthetic data of the MAIA dataset, where we used OpenAI’s GPT-4o API, the experiments on MAIA were conducted using A100100 GPUs (4040GB). Overall, the total computational budget was on the order of $\sim$1,000 GPU hours.
@@ B.1 Models tested
We benchmarked six recent VLMs. Both experiments and related quantitative evaluation have been done using lmms-eval (Zhang et al., 2024a), a framework for the evaluation of multimodal models.
InternVL2.
Chen et al. (2024): 88B parameter transformer-based multimodal model employing advanced cross-attention; pre-trained on large-scale image-text and video datasets for diverse multimodal tasks and instruction-tuned. It uses InternLM2.5 as open-sourced-77B parameter chat model. Hugging Face model: OpenGVLab/InternVL2-8B.
InternVL3.
Zhu et al. (2025): 7878B parameter multimodal model trained with a native multimodal pre-training paradigm, integrating linguistic and visual capabilities from the start. It employs Variable Visual Position Encoding (V2PE) and advanced post-training strategies, achieving state-of-the-art performance among open-source VLMs. Hugging Face model: OpenGVLab/InternVL3-78B.
LLaVA-NeXT-Video.
Zhang et al. (2024b): 77B parameter model built on the LLaVA framework, optimized for video understanding with mechanisms to capture temporal dynamics; fine-tuned on video instruction data. Base LLM: Vicuna-77B (v1.51.5). Hugging Face model: llava-hf/LLaVA-NeXT-Video-7B-hf.
LLaVa-OneVision.
Li et al. (2024a): 77B parameter model that builds on the LLaVA framework with a Qwen2 LLM backbone to serve as a general-purpose vision-language assistant; pre-trained on extensive multimodal data to deliver robust cross-modal reasoning. Hugging Face model: lmms-lab/llava-onevision-qwen2-7b-ov.
Qwen2.5-VL.
Qwen et al. (2025): 77B and 7272B parameter VLMs of the Qwen family using the Qwen2.52.5 LLM decoder; key enhancements are related to grounding, working with longer videos and capturing events. It was pre-trained on comprehensive visual and textual datasets and fine-tuned for detailed, context-aware responses. Hugging Face model: Qwen/Qwen2.5-VL-7B-Instruct and Qwen/Qwen2.5-VL-72B-Instruct.
@@
As described in Section 4we used five open-weight LLMs which have shown good performance on a variety of tasks on Italian Magnini et al. (2025). For conducting these experiments, we used Minicons library Misra (2022), a high-level wrapper around Hugging Face for investigating predictive behavior of transformer models. Specifically, probabilities were computed adding a normalization parameter to take into account the different length of sentences in terms of tokens. Models used in the Hugging face Hub are: Llama-$3.1$ (88B-Instruct), LLaMAntino-22 (77B), LLaMAntino-33-ANITA (88B-Instruct), Gemma (77B) and Qwen2.52.5 (77B-Instruct).
B.2 Tasks Details
Table 5 provide details with respect to the results obtained in Task 1 without considering any form of aggregation into pools (i.e. single-accuracy). Figures 6, 7, and 8 represent the fingerprint of models through MAIA’s reasoning categories. Specifically, Figure 6 reports results for Task 1, Figure 7 for Task 1 when models make 8/8 correct choices within the 8 TS-FS pairs that make up the pools, and Figure 8 for Task 2.
As regards the generation task, we combined similarity-based metrics with an LLM-as-a-judge approach, the latter being more suitable for handling open-ended responses. Using 8 reference answers for evaluating the generation correctness of VLMs allowed us to prioritize semantic alignment over surface similarity, following Lee et al. (2024a) and Mañas et al. (2024). GPT-4o was adopted as evaluation model, and its judgments showed high agreement with human annotators (Fleiss’ Kappa: 0.82). Figure 5 shows the prompt used for this evaluation.